OpenStudy (anonymous):

The area of a rectangle is 24 square meters. Find the length and width of the rectangle if its length is 5 meters greater than its width. Use an equation and the formula for Area of a rectangle = (width)(length).

OpenStudy (anonymous):

L xW =24 L=w+5 this means 24 = (w+5) x w 24= w^2 +5w w^2 +5w -24 = 0 (w+8)(w-3)=0 w = -8 or 3 w is positive W=3

OpenStudy (anonymous):

L = w+5 =3+5 +8